# Optimal Consumption and Portfolio Choice with No-Borrowing Constraint in the Kim-Omberg Model

### 03 Mar 2026 v1

@@ Appendix A Technical Proofs
[…]
@@
$d\hat{\beta}_{t}$ $=\sigma_{\beta}\,dW^{\beta,\mathbb{Q}}_{t}+\bigg(\kappa(\overline{\beta}-\hat{\beta}_{t})-\rho\,\frac{\hat{\beta}_{t}}{\sigma}\sigma_{\beta}\bigg)\,dt,\quad t>0,\quad\hat{\beta}_{0}=\beta\in\mathbb{R}.$
We set $\hat{\mathbb{F}}:=(\hat{\mathcal{F}}_{t})_{t}$, where $\hat{\mathcal{F}}_{t}:=\mathcal{F}_{t}^{W,\mathbb{Q}}\vee\mathcal{F}_{t}^{W^{\beta},\mathbb{Q}}$, $t\geq 0$. We then define the following value functions (where the infimum is taken over all $\hat{\mathbb{F}}-$stopping times)
@@
$\mathbb{E}^{\mathbb{Q}}_{z,\beta}\bigg[\int_{0}^{\infty}e^{-rt}\hat{Z}^{-\frac{1}{\gamma}}_{t}dt\bigg]=\int_{0}^{\infty}e^{-rt}\mathbb{E}^{\mathbb{Q}}_{z,\beta}[\hat{Z}_{t}^{-\frac{1}{\gamma}}]\,dt,$
and we recall that
$W_{t}^{\mathbb{Q}}=\rho\,W_{t}^{\beta,\mathbb{Q}}+\sqrt{1-\rho^{2}}\,\widehat{W}^{1,\perp}_{t},\quad t\geq 0,$
due to the correlation between $W^{\mathbb{Q}}$ and $W^{\beta,\mathbb{Q}}$, where $\widehat{W}^{1,\perp}$ is a Brownian motion independent of $W^{\beta,\mathbb{Q}}$. By using the explicit representation of the strong solution to (3.16), we find
@@
[…]
$=z^{-\frac{1}{\gamma}}\mathbb{E}^{\mathbb{Q}}_{1,\beta}\bigg[\mathbb{E}^{\mathbb{Q}}\bigg[\exp\bigg(-\frac{1}{\gamma}\int_{0}^{t}(\delta-r+\tfrac{1}{2}\tfrac{\hat{\beta}_{s}^{2}}{\sigma^{2}})ds$
$\quad+\frac{1}{\gamma}\rho\int_{0}^{t}\tfrac{\hat{\beta}_{s}}{\sigma}dW^{\beta,\mathbb{Q}}_{s}+\frac{1}{\gamma}\sqrt{1-\rho^{2}}\int_{0}^{t}\tfrac{\hat{\beta}_{s}}{\sigma}d\widehat{W}^{1,\perp}_{s}\bigg)\bigg|\mathcal{F}_{t}^{W^{\beta},\mathbb{Q}}\bigg]\bigg]$
$=z^{-\frac{1}{\gamma}}\mathbb{E}^{\mathbb{Q}}_{1,\beta}\bigg[\exp\bigg(-\frac{1}{\gamma}\int_{0}^{t}(\delta-r+\tfrac{1}{2}\tfrac{\hat{\beta}_{s}^{2}}{\sigma^{2}})ds+\frac{1}{\gamma}\rho\int_{0}^{t}\tfrac{\hat{\beta}_{s}}{\sigma}dW^{\beta,\mathbb{Q}}_{s}+\tfrac{1}{2}\frac{1}{\gamma^{2}}(1-\rho^{2})\int_{0}^{t}\tfrac{\hat{\beta}_{s}^{2}}{\sigma^{2}}ds\bigg)\bigg]$
[…]
@@
$\quad\cdot\mathbb{E}^{\mathbb{Q}}_{1,\beta}\bigg[\exp\bigg(\bigg(\tfrac{\frac{1}{\gamma^{2}}(1-\rho^{2})-\frac{1}{\gamma}}{2\sigma^{2}}\bigg)\int_{0}^{t}\hat{\beta}_{s}^{2}ds+\tfrac{\rho}{\sigma\gamma}\int_{0}^{t}\hat{\beta}_{s}dW^{\beta,\mathbb{Q}}_{s}-\tfrac{1}{2}\tfrac{\rho^{2}}{\sigma^{2}\gamma^{2}}\int_{0}^{t}\hat{\beta}_{s}^{2}ds+\tfrac{1}{2}\tfrac{\rho^{2}}{\sigma^{2}\gamma^{2}}\int_{0}^{t}\hat{\beta}_{s}^{2}ds\bigg)\bigg]$
$=z^{-\frac{1}{\gamma}}\exp\big(-\frac{1}{\gamma}(\delta-r)t\big)$
$\quad\cdot\mathbb{E}^{\mathbb{Q}}_{1,\beta}\bigg[\exp\bigg(\tfrac{1-\gamma}{2\sigma^{2}\gamma^{2}}\int_{0}^{t}\hat{\beta}_{s}^{2}ds\bigg)\exp\bigg(\tfrac{\rho}{\sigma\gamma}\int_{0}^{t}\hat{\beta}_{s}dW^{\beta,\mathbb{Q}}_{s}-\tfrac{1}{2}\tfrac{\rho^{2}}{\sigma^{2}\gamma^{2}}\int_{0}^{t}\hat{\beta}_{s}^{2}ds\bigg)\bigg].$
@@
$N_{t}:=\exp\bigg(\tfrac{\rho}{\sigma\gamma}\int_{0}^{t}\hat{\beta}_{s}dW^{\beta,\mathbb{Q}}_{s}-\tfrac{1}{2}\tfrac{\rho^{2}}{\sigma^{2}\gamma^{2}}\int_{0}^{t}\hat{\beta}_{s}^{2}ds\bigg)$
@@
$\frac{d\mathbb{Q^{\prime}}}{d\mathbb{Q}}\bigg|_{\hat{\mathcal{F}}_{t}}=N_{t},$
@@
$=\frac{1}{\gamma}\xi^{-\frac{1}{\gamma}-1}\,\varepsilon\,\mathbb{E}^{\mathbb{Q}}\Big[\int_{0}^{\infty}e^{-rt}(\hat{Z}^{1,\beta}_{t})^{-\frac{1}{\gamma}}dt\Big],$
@@
(A.8) $\hat{\beta}^{\beta}_{t}=\beta e^{-at}+b(1-e^{-at})+\sigma_{\beta}\int_{0}^{t}e^{-a(t-s)}dW_{s}^{\beta,\mathbb{Q}},$
@@
$1<p<\min\bigg\{\gamma,\frac{\gamma\sigma\bigg(\kappa+\frac{\rho\sigma_{\beta}}{\sigma}\bigg)}{|\rho|\sigma_{\beta}}\bigg\}$
@@
(A.14) $B_{t}$ $:=\frac{\sigma_{\beta}}{\sigma^{2}}\int_{0}^{t}e^{-as}\Big(\int_{0}^{s}e^{-a(s-u)}dW_{u}^{\beta,\mathbb{Q}}\Big)ds-\frac{1}{\sigma}\int_{0}^{t}e^{-as}dW_{s}^{\mathbb{Q}}.$
@@
$B_{t}$ $=\frac{\sigma_{\beta}}{\sigma^{2}}\int_{0}^{t}\left(\int_{u}^{t}e^{-as}e^{-a(s-u)}ds\right)dW_{u}^{\beta,\mathbb{Q}}-\frac{1}{\sigma}\int_{0}^{t}e^{-au}dW_{u}^{\mathbb{Q}}$
(A.16) $=\frac{\sigma_{\beta}}{2a\sigma^{2}}\int_{0}^{t}\left(e^{-au}-e^{-2at+au}\right)dW_{u}^{\beta,\mathbb{Q}}-\frac{1}{\sigma}\int_{0}^{t}e^{-au}dW_{u}^{\mathbb{Q}}$
$=\frac{\sigma_{\beta}\rho}{2a\sigma^{2}}\int_{0}^{t}\left(e^{-au}-e^{-2at+au}\right)dW_{u}^{\mathbb{Q}}+\frac{\sigma_{\beta}\sqrt{1-\rho^{2}}}{2a\sigma^{2}}\int_{0}^{t}\left(e^{-au}-e^{-2at+au}\right)dW_{u}^{2,\perp}$
$\quad-\frac{1}{\sigma}\int_{0}^{t}e^{-au}dW_{u}^{\mathbb{Q}}$
$=\int_{0}^{t}\underbrace{\left(\frac{\sigma_{\beta}\rho}{2a\sigma^{2}}(e^{-au}-e^{-2at+au})-\frac{1}{\sigma}e^{-au}\right)}_{=:f(u,t)}dW_{u}^{\mathbb{Q}}+\int_{0}^{t}\underbrace{\frac{\sigma_{\beta}\sqrt{1-\rho^{2}}}{2a\sigma^{2}}\left(e^{-au}-e^{-2at+au}\right)}_{=:g(u,t)}d\widehat{W}_{u}^{2,\perp},$
where we used $W_{u}^{\beta,\mathbb{Q}}=\rho W_{u}^{\mathbb{Q}}+\sqrt{1-\rho^{2}}\widehat{W}_{u}^{2,\perp}$ due to $Corr(W_{u}^{\mathbb{Q}},W_{u}^{\beta,\mathbb{Q}})=\rho u$. The deterministic functions $f(u,t)$ and $g(u,t)$ in (A.16) can be bounded as follows
$|f(u,t)|\leq e^{-au}\left(\frac{\sigma_{\beta}|\rho|}{2a\sigma^{2}}+\frac{1}{\sigma}\right)=:C_{f}e^{-au},\quad|g(u,t)|\leq\frac{\sigma_{\beta}\sqrt{1-\rho^{2}}}{2a\sigma^{2}}e^{-au}=:C_{g}e^{-au},$
and since $W^{\mathbb{Q}}$ is independent of $\widehat{W}^{2,\perp}$, we obtain
(A.17) $\mathrm{Var}(B_{t})$ $=\int_{0}^{t}f(u,t)^{2}du+\int_{0}^{t}g(u,t)^{2}du\leq\frac{C_{f}^{2}+C_{g}^{2}}{2a}=:K.$
@@
(A.22) $\bigg\{\left|(\hat{Z}_{t}^{\beta_{\varepsilon}})^{-1/\gamma}\left(\int_{0}^{t}e^{-as}\frac{\hat{\beta}_{s}^{\beta_{\varepsilon}}}{\sigma^{2}}ds-\frac{1}{\sigma}\int_{0}^{t}e^{-as}dW_{s}^{\mathbb{Q}}\right)\right|\mathbf{1}_{\{t\leq\tau^{*}\}}\bigg\}_{\varepsilon\in(0,\varepsilon_{0})}$
@@
$1<\tilde{p}<\min\bigg\{\gamma,\frac{\gamma\sigma\bigg(\kappa+\frac{\rho\sigma_{\beta}}{\sigma}\bigg)}{|\rho|\sigma_{\beta}}\bigg\},$
@@
$\sup_{\varepsilon\leq\varepsilon_{0}}\mathbb{E}^{\mathbb{Q}}\!\left[\left|(\hat{Z}_{t}^{z,\beta_{\varepsilon}})^{-\frac{1}{\gamma}}\left(\int_{0}^{t}e^{-as}\frac{\hat{\beta}_{s}^{\beta_{\varepsilon}}}{\sigma^{2}}ds-\frac{1}{\sigma}\int_{0}^{t}e^{-as}dW_{s}^{\mathbb{Q}}\right)\right|^{m}\mathbf{1}_{\{t\leq\tau^{*}\}}\right]\leq z^{-\frac{s}{\gamma}}\,e^{-\frac{s}{\gamma}(\delta-r)t}M_{mq}(\beta)^{m},$

### 11 Mar 2026 v3

@@ Appendix A Technical Proofs
Under $\mathbb$$\bar${P}}, we then have for any $\mathbb{F}^{W}$-stopping time $\tau$:
@@
$d\hat{\beta}_{t}$ $=-\sigma_{\beta}\,dW^{\mathbb{Q}}_{t}+\bigg(\kappa(\overline{\beta}-\hat{\beta}_{t})+\,\frac{\hat{\beta}_{t}}{\sigma}\sigma_{\beta}\bigg)\,dt,\quad t>0,\quad\hat{\beta}_{0}=\beta\in\mathbb{R}.$
We then define the following value functions (where the infimum is taken over all $\mathbb{F}^{W,\mathbb{Q}}$-stopping times)
@@
$\mathbb{E}^{\mathbb{Q}}_{z,\beta}\bigg[\int_{0}^{\infty}e^{-rt}\hat{Z}^{-\frac{1}{\gamma}}_{t}dt\bigg]=\int_{0}^{\infty}e^{-rt}\mathbb{E}^{\mathbb{Q}}_{z,\beta}[\hat{Z}_{t}^{-\frac{1}{\gamma}}]\,dt.$
By using the explicit representation of the strong solution to (3.16), we find
@@
$=z^{-\frac{1}{\gamma}}\mathbb{E}^{\mathbb{Q}}_{1,\beta}\bigg[\exp\bigg(-\frac{1}{\gamma}(\delta-r)t-\tfrac{1}{2\sigma^{2}\gamma}\int_{0}^{t}\hat{\beta}_{s}^{2}ds+\frac{1}{\gamma}\int_{0}^{t}\tfrac{\hat{\beta}_{s}}{\sigma}dW_{s}^{\mathbb{Q}}\bigg)\bigg]$
$=z^{-\frac{1}{\gamma}}\exp\big(-\frac{1}{\gamma}(\delta-r)t\big)\cdot$
$\quad\cdot\mathbb{E}^{\mathbb{Q}}_{1,\beta}\bigg[\exp\bigg(-\frac{1}{2\sigma^{2}\gamma}\int_{0}^{t}\hat{\beta}_{s}^{2}ds+\tfrac{1}{\sigma\gamma}\int_{0}^{t}\hat{\beta}_{s}dW^{\mathbb{Q}}_{s}-\tfrac{1}{2}\tfrac{1}{\sigma^{2}\gamma^{2}}\int_{0}^{t}\hat{\beta}_{s}^{2}ds+\tfrac{1}{2}\tfrac{1}{\sigma^{2}\gamma^{2}}\int_{0}^{t}\hat{\beta}_{s}^{2}ds\bigg)\bigg]$
@@
$\quad\cdot\mathbb{E}^{\mathbb{Q}}_{1,\beta}\bigg[\exp\bigg(\tfrac{1-\gamma}{2\sigma^{2}\gamma^{2}}\int_{0}^{t}\hat{\beta}_{s}^{2}ds\bigg)\exp\bigg(\tfrac{1}{\sigma\gamma}\int_{0}^{t}\hat{\beta}_{s}dW^{\mathbb{Q}}_{s}-\tfrac{1}{2}\tfrac{1}{\sigma^{2}\gamma^{2}}\int_{0}^{t}\hat{\beta}_{s}^{2}ds\bigg)\bigg].$
@@
$N_{t}:=\exp\bigg(\tfrac{1}{\sigma\gamma}\int_{0}^{t}\hat{\beta}_{s}dW^{\mathbb{Q}}_{s}-\tfrac{1}{2}\tfrac{1}{\sigma^{2}\gamma^{2}}\int_{0}^{t}\hat{\beta}_{s}^{2}ds\bigg)$
@@
$\frac{d\mathbb{Q^{\prime}}}{d\mathbb{Q}}\bigg|_{\mathcal{F}^{W,\mathbb{Q}}_{t}}=N_{t},$
@@
$=\frac{1}{\gamma}\zeta^{-\frac{1}{\gamma}-1}\,\varepsilon\,\mathbb{E}^{\mathbb{Q}}\Big[\int_{0}^{\infty}e^{-rt}(\hat{Z}^{1,\beta}_{t})^{-\frac{1}{\gamma}}dt\Big],$
@@
(A.8) $\hat{\beta}^{\beta}_{t}=\beta e^{-at}+b(1-e^{-at})-\sigma_{\beta}\int_{0}^{t}e^{-a(t-s)}dW_{s}^{\mathbb{Q}},$
@@
$1<p<\min\bigg\{\gamma,\frac{\gamma\sigma\bigg(\kappa-\frac{\sigma_{\beta}}{\sigma}\bigg)}{\sigma_{\beta}}\bigg\}$
@@
(A.14) $B_{t}$ $:=-\frac{\sigma_{\beta}}{\sigma^{2}}\int_{0}^{t}e^{-as}\Big(\int_{0}^{s}e^{-a(s-u)}dW_{u}^{\mathbb{Q}}\Big)ds-\frac{1}{\sigma}\int_{0}^{t}e^{-as}dW_{s}^{\mathbb{Q}}.$
@@
$B_{t}$ $=-\frac{\sigma_{\beta}}{\sigma^{2}}\int_{0}^{t}\left(\int_{u}^{t}e^{-as}e^{-a(s-u)}ds\right)dW_{u}^{\mathbb{Q}}-\frac{1}{\sigma}\int_{0}^{t}e^{-au}dW_{u}^{\mathbb{Q}}$
(A.16) $=-\frac{\sigma_{\beta}}{2a\sigma^{2}}\int_{0}^{t}\left(e^{-au}-e^{-2at+au}\right)dW_{u}^{\mathbb{Q}}-\frac{1}{\sigma}\int_{0}^{t}e^{-au}dW_{u}^{\mathbb{Q}}$
$=\int_{0}^{t}\underbrace{\left(-\frac{\sigma_{\beta}}{2a\sigma^{2}}(e^{-au}-e^{-2at+au})-\frac{1}{\sigma}e^{-au}\right)}_{=:f(u,t)}dW_{u}^{\mathbb{Q}}.$
The deterministic function $f(u,t)$ in (A.16) can be bounded as follows
$|f(u,t)|\leq e^{-au}\left(\frac{\sigma_{\beta}}{2a\sigma^{2}}+\frac{1}{\sigma}\right)=:C_{f}e^{-au},$
and hence, we obtain
(A.17) $\mathrm{Var}(B_{t})$ $=\int_{0}^{t}f(u,t)^{2}du\leq\frac{C_{f}^{2}}{2a}=:K.$
@@
(A.22) $\bigg\{\left|(\hat{Z}_{t}^{z,\beta_{\varepsilon}})^{-1/\gamma}\left(\int_{0}^{t}e^{-as}\frac{\hat{\beta}_{s}^{\beta_{\varepsilon}}}{\sigma^{2}}ds-\frac{1}{\sigma}\int_{0}^{t}e^{-as}dW_{s}^{\mathbb{Q}}\right)\right|\mathbf{1}_{\{t\leq\tau^{*}\}}\bigg\}_{\varepsilon\in(0,\varepsilon_{0})}$
@@
$1<\tilde{p}<\min\bigg\{\gamma,\frac{\gamma\sigma\bigg(\kappa-\frac{\sigma_{\beta}}{\sigma}\bigg)}{\sigma_{\beta}}\bigg\},$
@@
$\sup_{\varepsilon\leq\varepsilon_{0}}\mathbb{E}^{\mathbb{Q}}\!\left[\left|(\hat{Z}_{t}^{z,\beta_{\varepsilon}})^{-\frac{1}{\gamma}}\left(\int_{0}^{t}e^{-as}\frac{\hat{\beta}_{s}^{\beta_{\varepsilon}}}{\sigma^{2}}ds-\frac{1}{\sigma}\int_{0}^{t}e^{-as}dW_{s}^{\mathbb{Q}}\right)\right|^{m}\mathbf{1}_{\{t\leq\tau^{*}\}}\right]\leq z^{-\frac{m}{\gamma}}\,e^{-\frac{m}{\gamma}(\delta-r)t}M_{mq}(\beta)^{m},$